Design and draw pie type high pass filter having cut of frequency of 2kHz and load resistance of 500Ω
1
Expert's answer
2019-03-25T09:08:13-0400

Design a first-order analog passive high-pass filter with the required properties. And since its cutoff frequency is 2 kHz, find its capacitance:


"f_c=\frac{1}{2\pi CR},"

"C=\frac{1}{2\pi f_cR}=1.592\cdot 10^{-7}\text{ F, or 159.2 nF}."